# FlashGS: Efficient 3D Gaussian Splatting for Large-scale and High-resolution Rendering

### v1

@@ Abstract.
This work introduces FlashGS, an open-source CUDA Python library  [^1] Available at https://github.com/InternLandMark/FlashGS. designed to facilitate the efficient differentiable rasterization of 3D Gaussian Splatting through algorithmic and kernel-level optimizations. FlashGS is developed based on the observations from a comprehensive analysis of rendering process to enhance computational efficiency and bring the technique to wide adoption. The paper includes a suite of optimization strategies, encompassing redundancy elimination, efficient pipelining, refined control and scheduling mechanisms, and memory access optimizations, all of which are meticulously integrated to amplify the performance of the rasterization process. An extensive evaluation of FlashGS’ performance has been conducted across a diverse spectrum of synthetic and real-world large-scale scenes, encompassing a variety of image resolutions. The empirical findings demonstrate that FlashGS consistently achieves an average 4x acceleration over mobile consumer GPUs, coupled with reduced memory consumption. These results underscore the superior performance and resource optimization capabilities of FlashGS, positioning it as a formidable tool in the domain of 3D rendering..
3D Gaussian Splatting, CUDA optimization, large-scale rendering
@@ 6.1.1. Testbed
We conduct our experiments on two NVIDIA GPU platforms. Most tests are performed on a consumer-grade RTX 3090 (Ampere architecture, 24GB GDDR6X memory, CUDA Compute Capability 8.6). We use an A100 (Ampere architecture, 80GB HBM2e, CUDA Compute Capability 8.0) data-center GPU to study performance sentivity. We compile our codes with GCC 10.5.0 and NVCC in CUDA 12.0.
@@ 6.1.2. Dataset
We comprehensively test FlashGS and 3DGS on 6 representative datasets, which cover very different resolutions and scenes. The Truck and Train are outdoor scenes from the Tanks & Temple. The Playroom and DrJohnson are two indoor scenes in DeepBlending. 3DGS claims that they achieve real-time rendering with 1080p resolution on these datasets. We use another 2 large-scale and high-resolution datasets which beyond 3DGS’s capabilities to render in real-time. MatrixCity (Li et al., 2023) is a comprehensive and large-scale high-quality dataset. Utilizing Unreal Engine 5, this dataset features two virtual cities covering a total area of 28km2, comprising 67,000 aerial images and 452,000 street images at a resolution of 1080p. Rubble in Mill19 contains thousands of high-resolution images captured by drones. We train 30K iterations on each dataset with 3DGS to obtain the Gaussian model for rendering.
@@ 6.1.3. Baseline
We compare FlashGS with 3DGS (Kerbl et al., 2023) algorithm and the some recent optimizations, gsplat (Ye and Kanazawa, 2023) and GScore (Lee et al., 2024). Since GScore is implemented as a hardware accelerator unit, we cannot directly test it. Its experimental results are derived from the speedup descriptions of its algorithmic optimizations within GScore  [^2] They claim that their shape-aware intersection test provides 1.71$\times$ speedup and subtle skipping offers an additional 15% improvement..
@@ 6.2. Overall Performance
As shown in Table 1, FlashGS outperforms the original 3DGS algorithm across all the performance metrics on every dataset. We can always achieve ¿ 100fps rendering on 3090, even for high-resolution and large-scale datasets. In the slowest frame in Rubble of all datasets, we achieve 124.8 fps. This demonstrate that FlashGS can perform real-time rendering even in extreme large and high-resolution cases. Flash GS achieves up to 25.39$\times$ speedup with an average 12.18$\times$ on the Matrixcity dataset at 4k resolution. We achieve 6.4$\times$ average speedup on all 11 scenes, while achieve 8.7$\times$ speedup on the 5 large-scale or high-resolution tests. FlashGS achieves 1.1 to 25.39 speedup for all the frames in different scenes, comprehensively proving the generality of our design and optimization.
@@ 6.3. Runtime Breakdown Analysis
Figure 13. Breakdown of the running time.
In this section, we compare and analyze the runtime breakdown of FlashGS to 3DGS, aiming to reveal the source of our excellent performance. Figure 13 shows the rasterizer time and breakdown for several frames of each scene, demonstrating that we have accelerated all stages, including preprocessing, sorting, and rendering.

### v2

@@ Abstract.
This work introduces FlashGS, an open-source CUDA Python library  [^1] Available at https://github.com/InternLandMark/FlashGS. designed to facilitate the efficient differentiable rasterization of 3D Gaussian Splatting through algorithmic and kernel-level optimizations. FlashGS is developed based on the observations from a comprehensive analysis of rendering process to enhance computational efficiency and bring the technique to wide adoption. The paper includes a suite of optimization strategies, encompassing redundancy elimination, efficient pipelining, refined control and scheduling mechanisms, and memory access optimizations, all of which are meticulously integrated to amplify the performance of the rasterization process. An extensive evaluation of FlashGS’ performance has been conducted across a diverse spectrum of synthetic and real-world large-scale scenes, encompassing a variety of image resolutions. The empirical findings demonstrate that FlashGS consistently achieves an average 4x acceleration over mobile consumer GPUs, coupled with reduced memory consumption. These results underscore the superior performance and resource optimization capabilities of FlashGS, positioning it as a formidable tool in the domain of 3D rendering.
3D Gaussian Splatting, CUDA optimization, Large-scale real-time rendering
@@ 6.1.1. Testbed
We conduct our experiments on two NVIDIA GPU platforms. Most tests are performed on a consumer-grade RTX 3090 (Ampere architecture, 24GB GDDR6X memory, CUDA Compute Capability 8.6). We use an A100 (Ampere architecture, 80GB HBM2e, CUDA Compute Capability 8.0) data-center GPU to study performance sensitivity. We compile our codes with GCC 10.5.0 and NVCC in CUDA 12.0. We have also integrated FlashGS as a callable module in Python to enhance usability.
@@ 6.1.2. Dataset
We comprehensively test FlashGS and 3DGS on 6 representative datasets, which cover very different resolutions on a total of 11 scenes with hundreds or thousands of frames per scene. The Truck and Train are outdoor scenes from the Tanks & Temple (Knapitsch et al., 2017). The Playroom and DrJohnson are two indoor scenes in DeepBlending (Hedman et al., 2018). 3DGS claims that they achieve real-time rendering with 1080p resolution on these datasets. We use another 2 large-scale and high-resolution datasets which beyond 3DGS’s capabilities to render in real-time. MatrixCity (Li et al., 2023) is a comprehensive and large-scale high-quality dataset, comprising thousands of aerial images at a resolution of 1080p. Rubble in Mill19 (Turki et al., 2022) contains high-resolution images captured by drones. We train 30K iterations on each dataset with 3DGS to obtain the Gaussian model for rendering.
@@ 6.1.3. Baseline
We compare FlashGS with 3DGS (Kerbl et al., 2023) and the recent optimizations on GScore (Lee et al., 2024). Since GScore is implemented as a hardware accelerator unit, we cannot directly test it. They claim that GScore achieves 1.86$\times$ speedup with their intersection and scheduling techniques  [^2] Their shape-aware intersection test provides 1.71$\times$ speedup and subtle skipping offers an additional 15% improvement..
@@ 6.2. Overall Performance
As shown in Table 1, the rasterization pipe of FlashGS outperforms the original 3DGS algorithm across all the performance metrics on every dataset with different scenes and resolutions. We can always achieve ¿ 100FPS rendering on RTX 3090, even for high-resolution and large-scale datasets. In the slowest frame in Rubble of all datasets, we achieve 107.3 FPS. This demonstrates that FlashGS can perform real-time rendering even in extremely large and high-resolution cases. Flash GS achieves up to 30.53$\times$ speedup with an average of 12.18$\times$ on the Matrixcity dataset at 4k resolution. We achieve 7.2$\times$ average speedup on all 11 scenes, while achieving 8.6$\times$ speedup on the 7 large-scale or high-resolution tests, which is 3.87$\times$ of the GScores results (Lee et al., 2024). FlashGS achieves a 2.29$\times$ to 30.53$\times$ speedup for all the frames in different scenes, comprehensively proving the generality of our design and optimization.
@@ 6.3. Runtime Breakdown Analysis
Figure 13. Rasterization runtime breakdown on 6 representative frames from different datasets, normalized to 3DGS.
In this section, we compare and analyze the runtime breakdown of FlashGS to 3DGS, aiming to reveal the source of our excellent performance. Figure 13 shows the rasterization time and the breakdown for 6 representative frames shows max or min speedup. This demonstrates that we have accelerated all stages, including preprocess, sort, and render. In FlashGS, these stages respectively account for average 19.6%, 29.1%, and 47.6% of the total time, whereas in 3DGS, they account for 13.2%, 25.4%, and 59.6%. The optimizations in the preprocess and render stages have already been discussed earlier, while the speedup in the sorting stage is primarily due to the reduction in the number of key-value pairs to be sorted after applying our precise intersection algorithm. We will give a more detailed analysis below.
6.3.1. Profiling Results
Figure 14. Profiling results of FlashGS: The number of instructions issued in rendering and the memory transactions in preprocessing. All results are normalized to 3DGS.
We further demonstrate our optimizations in the rendering and preprocessing stages through profiling results of memory and compute units, as shown in the figures. Figure  14 shows that we reduce the issued instructions in the rendering stage, alleviating the computation-bound problem. This prove the effectiveness of our precise intersection algorithm and the optimizations for low-latency rendering. The total issues instructions is significantly reduced by one to two orders of magnitude. For memory access trasactions in preprocessing, we also reduced the number of global memory accesses by 43%-87% compared to 3DGS.
Figure 15. Number of rendered Gaussian-tile (kv) pairs and instructions issued per pair of FlashGS (Normalized to 3DGS).
Figure 15 further shows the reason for the significant instruction reduction in the rendering stage, which is dominant in the rasterization. In tile-based rendering, the total number of issued instructions is the product of the number of tiles and the instructions per tile. Therefore, this reduction mainly stems from two aspects. In Figure 15, the number of rendered key-value pairs benefit from our intersection optimization, reducing by 68%-96%. The instructions involved in each tile’s computation also decrease by 67%-71%, as we further optimize the renderCUDA kernel at the instruction level. The reduction in the number of generated key-value pairs also benefits the sorting process, as it significantly decreases the size of the list to be sorted. Additionally, less memory is required to store these key-value pairs.
6.4. Sensitivity Study
In this section, we study the FlashGS performance sensitivity on different GPUs and sciences.
6.4.1. Performance on A100
We also conduct the experiments presented in Table 1 on the A100 GPU, and the results are shown in Table 1. The total rasterizerization time is slower on the A100 compared to the 3090, with on average 1.43$\times$ slower. This is primarily due to the rendering step is dominant, as shown in Figure 13, which heavily relies on FP32 computation, where the FP32 peak performance of the A100 is only 19.5TFLOPS while the 3090 is 35.6TFLOPS. Although the operations before rendering, such as sorting, are mainly memory-bandwidth-bound, and the A100 has a higher bandwidth than 3090, these benefits can not outweigh the render performance reduction. However, we still achieve significant speedups to 3DGS in all datasets on the A100 with an average of 123.8-423.7 FPS across 11 scenes in Table 1, always reaching real-time rendering.
6.4.2. Performance on different scenes
6.5. Image Quality
Figure 16. The rendering quality and rasterization time on Matricity(1080p)-800 frame.
Figure 17 shows the most complex frame of the origin large-scale MatrixCity dataset at 1080p, to compare the output quality. We compare the Peak Signal to Noise Ratio (PSNR) between FlashGS and 3DGS, which is a standard metric in computer graphics (higher is better). The result shows that FlashGS does not change the quality, keeping 31.52 PSNR with 8.57$\times$ speedup. It is reasonable that our precise intersection algorithms only reduce the false-positive redundancies. And we do not apply pruning or quantization strategies in our implementation so there is no accuracy loss.